OpenStudy (adorkable_bree):

use [1 -2 -3 7] encode the phrase "ONE QUESTION TO GO" 1 = A 2 = B 3 = C 4 = D 5 = E 6 = F 7 = G 8 = H 9 = I 10 = J 11 = K 12 = L 13 = M 14 = N 15 = O 16 = P 17 = Q 18 = R 19 = S 20 = T 21 = U 22 = V 23 = W 24 = X 25 = Y 26 = Z The matrix c= [1 -2 -3 7] was used to encode a phrase to [7 -28 -25 -35 -2 -21 107 90 123 17] Find c^-1 and use it to decode the matrix. PLEASE HELP I DONT UNDERSTAND THIS AT ALL

OpenStudy (freckles):

So I guess c is a 2 by 2 matrix or a 1 by 4 matrix?

OpenStudy (adorkable_bree):

2 by 2

OpenStudy (freckles):

So you can find the inverse of c?

OpenStudy (freckles):

\[(\left[\begin{matrix}a & b \\ c & d\end{matrix}\right])^{-1}=\frac{1}{ad-bc} \left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right]\]

OpenStudy (adorkable_bree):

give me a min

OpenStudy (adorkable_bree):

1/1 [ 7 -3 -2 1] ??

OpenStudy (adorkable_bree):

@freckles

OpenStudy (freckles):

It looks like you didn't change the sign of -3 and -2 but you switch their places...

OpenStudy (adorkable_bree):

wait so they would be positive...did i even do that right sorta?

OpenStudy (freckles):

\[(\left[\begin{matrix}a & b \\ c & d\end{matrix}\right])^{-1}=\frac{1}{ad-bc} \left[\begin{matrix}d & -b \\ -c & a\end{matrix}\right] \] \[ (\left[\begin{matrix}1 & -2 \\ -3 & 7\end{matrix}\right])^{-1}=\frac{1}{(1)(7)-(-2)(-3)} \left[\begin{matrix}7 & -(-2) \\ -(-3) & 1\end{matrix}\right] \]

OpenStudy (freckles):

I will let you simplify that.

OpenStudy (freckles):

When you are done I'm assuming that one matrix like in your last sentence is a 2 by 5 matrix So you are going to that your inverse of c and multiply the 2 by 5 matrix

OpenStudy (freckles):

This will give us what we need to decode the last matrix

OpenStudy (freckles):

We will encode the "ONE QUESTION TO GO" when we are done decrpying the 2 by 5 matrix

OpenStudy (adorkable_bree):

alright. is this an easier way of figuring it out ?

OpenStudy (freckles):

? Aren't we suppose to decode that last matrix in your last sentence?

OpenStudy (adorkable_bree):

yes

OpenStudy (freckles):

Ok so we have to find the inverse of c have you done that yet?

OpenStudy (freckles):

I gave you the setup and just asked you to simplify

OpenStudy (freckles):

You are going to multiply inverse of c times that 2 by 5 matrix

OpenStudy (freckles):

http://aix1.uottawa.ca/~jkhoury/cryptography.htm here is an example of what I interpreted that we are doing

OpenStudy (freckles):

They used a 3 by 3 key instead of a 2 by 2 key

OpenStudy (freckles):

But to decode something you do need the matrix's inverse whether it be a 3 by 3 or a 2 by 2 key

OpenStudy (adorkable_bree):

well that seems to explain it really well

OpenStudy (adorkable_bree):

yes it is

OpenStudy (freckles):

So after you perform that multiplication You will use that list list of letters you have above to translate

OpenStudy (freckles):

The only think is the message makes no sense

OpenStudy (freckles):

thing*

OpenStudy (adorkable_bree):

i think i was thinking to hard on these two problems and got all confused